OpenStudy (anonymous):

What is the slope of the line that passes through (2, 5) and (–1, 5)? –3 0 undefined 3

OpenStudy (anonymous):

i got 1/-3

OpenStudy (anonymous):

but i'm not sure what the answer would be

OpenStudy (rmrjr22):

y2-y1/x2-x1

OpenStudy (rmrjr22):

= m (slope)

OpenStudy (anonymous):

i got 1/-3 as my slope

OpenStudy (rmrjr22):

(5-5)/(-1-2) = 0/-3

OpenStudy (anonymous):

y coordinates are the same, so the line is horizontal and the slope is zero

OpenStudy (anonymous):

oh crap aahaha

OpenStudy (anonymous):

wait i dont understand

OpenStudy (anonymous):

the y coordinates are both 5

OpenStudy (anonymous):

i know the slope is 0/-3 but how do i figure out if it's undefined or not

OpenStudy (rmrjr22):

they cancel out wouldnt it be undefined?

OpenStudy (anonymous):



OpenStudy (rmrjr22):

ah, gotcha...

OpenStudy (anonymous):

the slope measures the steepness of the line but in this case the line is horizontal, as \(x\) increases by 1 unit, y stays the same, so the slope is zero over one, don't go up or down at all

OpenStudy (rmrjr22):

rise/run = slope

OpenStudy (rmrjr22):

there is no rise, but a run

OpenStudy (lopezking1):

(2, 5) and (–1, 5) gradient of line or slope : (y2 - y1) / (x2 - x1) (2, 5) let x1 represent 2 let y1 represent 5 (-1, 5) let x2 represent -1 let y2 represent 5 now using equation: (5 - 5) / (-1 - 2) = 0/-3

OpenStudy (rmrjr22):

so its 0

OpenStudy (anonymous):

@pamelasaucedo \(\frac{0}{\text{some number}}\) is \(0\)

OpenStudy (anonymous):

so is it 0 or undefined? you guys are confusing me lol

OpenStudy (rmrjr22):

0

OpenStudy (lopezking1):

gradient is 0. therefore the line is parallel to the x axis

OpenStudy (anonymous):

a fraction with zero in the numerator (top) is zero

OpenStudy (anonymous):

oh alright i get it now

OpenStudy (anonymous):

a fraction with zero in the denominator is undefined